### v1

# Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals

Abstract
Here we present the discovery and characterization of TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 family. We previously reported on the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) kagome metals. This structure type represents a magnetically diverse and exfoliable family of kagome metals. Detailed characterization of the magnetic and electronic transport properties of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state. The system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (-50%) near a metamagnetic transition. Unlike the Lu-Gd analogs, TbTi3Bi4 cannot be synthesized by the methodology presented previously due to phase competition with another, unreported class of materials, the Ln2-xTi6+xBi9 (x1.7–1.2) (Ln: Tb–Lu) shurikagome metals. Like the LnTi3Bi4 kagome metals, the Ln2-xTi6+xBi9 family exhibits quasi-2D networks of titanium and chains of rare-earth. We present the structures and some select basic properties of the Ln2-xTi6+xBi9 family alongside the newly discovered TbTi3Bi4. A discussion regarding the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families helps frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and provides a strategy to circumvent formation of Ln2-xTi6+xBi9. 111Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
I Introduction
Spurred on by continued discoveries in the AM3X5, AM6X6, and CoSn-type compounds, kagome metals have proven to be a rich space for exploring the intersection of solid-state chemistry and fundamental condensed matter physics. The impetus for much of the inspiration comes from the basic electronic structure afforded by the kagome motif – one which can contain Dirac points, flat bands, and Van Hove singularities. [1, 2, 3, 4]. Derived from the triangular lattice, the kagome motif have shown remarkable structural and chemical diversity, allowing for a rich interaction between structural chemistry, electronic properties, and magnetism. As a prime example, the AM3X5 compounds have been heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, superconductivity).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity.[14] In fact, embedding magnetism into the kagome motif underpins much of the research into the AM6X6 and CoSn-type compounds, which continues to provide the motivation for the development of new kagome materials.[15, 16, 17, 18, 16, 19, 20, 21, 22, 23, 24, 25, 26]
One fundamental way to develop new materials is to perturb simpler, “fundamental” prototype structures. Take, for example, the CoSn family of compounds. By adding a filler atom in the natural voids formed by the structure, the CoSn-type stacking gives way to another of the most researched kagome families, the AM6X6. Further perturbations on the AM6X6 prototype yield a grandchild of the CoSn prototype – the AM3X4 kagome metals.
We reported previously on the discovery of both the antimonide AV3Sb4 (A: Eu2+, Yb2+) [27] and bismuthide ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[28] which make up a larger family of AM3X4 compounds.[27, 28, 29, 30, 31, 32, 33] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals. The A-site elements comprise zig-zag chains which run parallel to the M-site kagome layers. Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds maintain a much more quasi-2D crystal structure. Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure. Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission (ARPES).[32, 34, 35, 36, 37, 38]
Under the conditions presented in our previous work, LnTi3Bi4 compounds containing rare-earth atoms smaller than Gd3+ did not appear. In this work, we present a modified growth procedure which allows stabilization of TbTi3Bi4, the smallest rare-earth cation reported in the LnTi3Bi4 structure thus far. Our investigation into the smaller rare-earth species (Tb–Lu) also unveiled an intimate competition between the LnTi3Bi4 kagome metals and a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals. By understanding the phase diagram surrounding the LnTi3Bi4 compound, we are able to subvert the formation of Tb2-xTi6+xBi9 and stabilize single crystals of TbTi3Bi4. This work details our rationale for the stabilization of TbTi3Bi4 alongside a discussion of its magnetic and electronic transport properties. Remarkably, TbTi3Bi4 presents some of the most rich and complex magnetism in the known AM3X4 compounds, reinforcing the need to understand and control the formation of new phases. Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals, which have strong chemical disorder and unusual magnetic behavior.
II Methods
II.1 Single Crystal Synthesis
TbTi3Bi4 single crystals are grown through a bismuth self-flux. Tb metal (Ames), Ti (Alfa 99.9% powder), and Bi (Alfa 99.999% low-oxide shot) were measured in a 2:3:20 ratio into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit.[39] The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules. The flux was heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h. The sample was then quenched by removing it from the furnace and immediately placing in a preheated 600°C furnace. The growth proceeded with a cooling rate of 2°C/hr to 400°C, where excess bismuth was removed through centrifugation. Ln2-xTi6+xBi9 single crystals are grown in a similar manner, except the samples are cooled linearly from 1100°C to 600°C at a rate of 2°C/hr before centrifugation at 600°C.
II.2 Bulk Characterization
Single crystals were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD). Diffraction data were collected at 100 K on a Bruker D8 Advance Quest diffractometer with a graphite monochromator using Mo K radiation ( = 0.71073 Å). Data integration, reduction, and structure solution was performed using the Bruker APEX4 software package. As noted previously, diffraction on LnTi3Bi4 crystals is challenging and requires face-indexing absorption corrections.[28] All atoms were refined with anisotropic thermal parameters. CIF files for all structures are included in the ESI. Laue diffraction utilizing a Photonic Science vertical Laue system was used to orient crystals of TbTi3Bi4 for physical property measurements.
Magnetization measurements of LnTi3Bi4 single crystals were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode. Additional measurements below 1.8 K on the MPMS3 utilized an additional He3 insert. Angle-resolved magnetization measurements were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
Heat capacity measurements were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS) equipped a dilution refrigerator insert. Thermal contact was achieved using Apezion N-grease. Electronic resistivity measurements were performed in a Quantum Design 9 T Dynacool PPMS. Single crystals were mounted to a sheet of sapphire and then the sapphire plate was then adhered to the sample puck stage using GE varish. The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane. Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%).
III Results and Discussion
III.1 Crystal Structure and Synthesis
Our prior work identified that the ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[28] However, growth of ATi3Bi4 crystals containing smaller rare-earth metals (Tb–Lu) was stymied due to preferential formation of Ln2-xTi6+xBi9. Figure 1(a) provides a simple schematic showing the lanthanide row sorted in order of the 9-coordinate Shannon ionic radius. Within the ATi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state. As such, the radii of Eu2+ and Yb2+ are analogous to the larger La3+ and Pr3+, respectively. Large cations preferentially stabilize the ATi3Bi4 phase (denoted by blue), and small cations preferentially stabilize Ln2-xTi6+xBi9 (green). Within this work, we demonstrate that Tb represents a unique cross-over between the two phases, and both TbTi3Bi4 and Tb2-xTi6+xBi9 can be obtained by tuning the growth profiles.
Figure 1(b,c) details the crystal structure of both the ATi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds. As we have discussed ATi3Bi4 extensively before, for simplicity we opt to show only half the unit cell here (the omitted half is generated through an inversion operation, space group Fmmm. ATi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln-Ln zig-zag chains. The local coordination of the Ln species is approximately 9-fold. Crystals of ATi3Bi4 are quasi-2D and can be easily exfoliated using adhesive tape. We have observed that both M-Sb and Ln-Sb terminations can be achieved.[27] The kagome network in ATi3Bi4 compounds is only slightly distorted, though the quasi-1D zig-zag chain limits the symmetry of the system to orthorhombic.
The Ln2-xTi6+xBi9 compounds can be understood as a heavily alloyed form of Ti8Bi9[40], where Ln atoms preferentially occupy two of the Ti sites. If the substitution was perfect, (e.g. Ln2Ti6Bi9) the Ln atoms would create a nearly ideal 1D chain that cuts perpendicular to the Ti–Ti network. The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to ATi3Bi4 (4Å)The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles. Paying homage to both the linguistic origin and coordination of the “kagome lattice,” the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
The two sites that are preferentially occupied by Ln atoms possess considerably different coordination than the remaining Ti sites in the parent Ti8Bi9 structure. Considering bonds 3.2Å, Ti atoms within the shurikagome network have an approximate coordination of 10 (4 Ti–Ti bonds, 6 Ti–Bi bonds) and a polyhedral volume of approximately 50.7Å3. Sites preferentially occupied by Ln possess a coordination of 8, solely from Ln–Bi bonds, and form slightly distorted square anti-prisms. The two mixed Ln/Ti sites are also distinct from one another, and one has a substantially larger polyhedral volume (56.9Å3) than the other (49.5Å3). As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and further preferentially occupy the largest of the mixed Ln/Ti sites. Generally speaking, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site. As such, the total stoichiometry can be more precisely written as Ln2-xTi6+xBi9 (x1.7–1.2). CIF files for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI. We will discuss trends in the Ln substitution alongside the magnetic properties in a later section.
As alluded to before, TbTi3Bi4 is not stable under conditions presented in our first manuscript.[28] Samples of large rare-earth ATi3Bi4 compounds are readily grown through a simple flux (cooled from 1100°C to 600°C at a rate of 2°C/hr). However, samples containing Tb will exclusively form Tb2-xTi6+xBi9. Figure 2(a) shows a small section of the Tb–Ti–Bi phase diagram near several of the phases of interest. Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[41, 42] Note the proximity of TbTi3Bi4 and Tb2-xTi6+xBi9 (particularly the proximity of the the hypothetically ideal Tb2Ti6Bi9). The large single-phase region of Ti8Bi9 (and the associated alloys with Tb2-xTi6+xBi9) is problematic, as it isolates a large swath of the phase diagram from Bi-rich fluxes. The single phase region even persists beyond where a hypothetical Bi2Ti6Bi9 exists (complete substitution of Bi on the mixed sites). There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[43] However, if Tb2Ti6Bi9 is not stable, then TbTi3Sb4 should still be accessible from a Bi self-flux.
We designed a series of experiments to test both the thermal stability of Tb2-xTi6+xBi9 and compositional dependence. We first observed that samples allowed to completely freeze (T=270°C) often exhibited Tb2-xTi6+xBi9 crystals that were rounded, as though they started to decompose. Modifying the flux towards Tb-rich conditions instead resulted in rounded crystals of TbBi. In both cases, the corners of the decomposing crystals were encased in agglomerates of TbTi3Bi4. This highlighted that TbTi3Bi4 is a low-temperature phase. Subsequent quenching experiments further suggested that TbTi3Bi4 is likely only stable below 600°C. Samples grown above 600°C reveal two alkemade triangles; 1) three phase TbBi–Tb2-xTi6+xBi9–Bi and 2) two phase Tb2-xTi6+xBi9–Bi. Samples grown below 600°C reveal instead; 1) three phase TbBi–TbTi3Bi4–Bi and 2) three phase TbTi3Bi4–Tb2-xTi6+xBi9–Bi. Figure 2(b) presents a simple projection of Figure 2(a) along the TiBi–TbBi pseudobinary line, summarizing the relative proximity and temperature stability of these results.
Crystals of TbTi3Sb4 are 1-5 mm pseudo-hexagonal crystals with a brilliant metallic luster. They are readily exfoliated and are lightly air-sensitive. Samples can be worked with in air, though they will swell and spall under extended (1 day) exposure. Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster. Samples are readily cleaved with a blade parallel to the ab-plane. Ln2-xTi6+xBi9 crystals are substantially more sensitive to air, and crystals will completely disintegrate within a few hours. In both cases, the samples expand and self-cleave along the ab-plane upon extended exposure to air.
III.2 Properties of TbTi3Bi4
Due to the reduced symmetry (Fmmm), the ATi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains. TbTi3Bi4 is no exception. To diagnose the potential anisotropies in the material, we first begin with a study of the magnetization along different principle crystallograpic axes. Figure 3(a) presents the magnetization of a single crystal of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two primary transitions at 24 K and 3 K. The [100] direction corresponds to the “easy axis,” and orienting the magnetic field parallel to the [010] or [001] induces a substantially weaker effect. The behavior is reminiscent of GdTi3Bi4[28], though crystals of TbTi3Bi4 respond negligibly to fields oriented along [001].
Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. Upon cursory inspection, the data again resembled that of GdTi3Bi4 (two metamagnetic transitions when ). However, the metamagnetic transitions in TbTi3Bi4 exhibit substantial field hysteresis. The data saturates at approximately The first transition plateaus at approximately 9.2 , in good agreement with the 9 expected for Tb3+. The first metamagnetic transition first plateaus at a moment of 2.9-3.2 , which is approximately 1/3 of the expected net moment. Furthermore, closer inspection at the second metamagnetic transition reveals an unusual inflection midway through the rising field direction (see Figure 3(b,inset)). This inflection occurs at approximately 6.2 slightly above 2/3 the expected moment. This inflection and other properties of the isothermal magnetization curves will be studied in detail below. Similar to the temperature-dependent magnetization, there are only weak, linear responses when or . Figure 3(c) provides a series of magnetization curves collected while rotating between , , and in 15°increments. The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further easy-[100]) anisotropy.
Figure 3(c) presents the inverse susceptibility, collected under a 100 Oe field where . The high temperature data exhibits a slight upwards curvature, potentially due to a temperature-independent contribution from core diamagnetism (bismuth) or the quartz sample holder. We show Curie-Weiss fits to the data with (gray) and without (black) a temperature-independent term. While the gray curve fits over a longer range of data, it requires a large emu mol-1 Oe-1 and yields an unphysical effective moment of 11 . We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit, which yields a =53 K and a = 9.9. The effective moment agrees well with that expected of Tb3+ (9.72), though the Weiss temperature is unphysically large (53 K), likely affected by the temperature independent contribution, limited temperature range, and crystal-field effects.
As a consistency check with the magnetization data, we performed zero-field heat capacity measurements down to 100 mK. The sample was initially oriented with such that measurements could also be made for the intermediate plateaus seen in Figure 2(b). The first order hysteretic nature of the transitions seen in the isothermal magnetization made the quantative analysis of field-dependent heat capacity challenging, and we have opted to reserve this data for a later study. Instead, we focus on the analysis of the zero-field heat capacity utilizing samples of LaTi3Bi4 as a nonmagnetic standard. Figure 3(d) demonstrates the experimental heat capacity data (green) as a function of temperature, plotted in . The two magnetic transitions are evident around 20 K and 2 K. At temperatures below 2 K, a nuclear Schottky peak dominates the heat capacity signal. The peak is still increasing at 0.2 K, which agrees relatively well with a Tb-based Schottky peak ( K).[44] Utilizing LaTi3Bi4 as the nonmagnetic standard (black) and invoking a simple nuclear Schottky model , we can extract the magnetic entropy for TbTi3Bi4.
Figure 3(e) shows the integrated magnetic entropy (green) as a function of temperature. The extracted magnetic contribution to the heat capacity () is also shown (gray) as a schematic reference for the magnetic transition temperatures (not to scale). Interestingly, integrating up to the primary magnetic transition appears to correspond with an entropy release of approximately . Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other ATi3Bi4 materials.[28] By 100 K we recover approximately 90% of the total entropy expected for Tb3+. Considering that the higher temerature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
We now return to the hysteretic metamagnetic transitions observed in Figure 3(b). Closer inspection into the second metamagnetic transition reveals a complex magnetic landscape which can only be elucidated with high density phase maps. Figure 4(a,d) present our efforts to map the magnetic landscape of TbTi3Bi4, demonstrating a set of isothermal magnetization curves collected in both the increasing (a) and decreasing (d) field directions. One can immediately see that the sharpness of the second transition belies a network of subtle changes in the magnetization. Taking the derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e). Immediately one can see that the plateau visible in the 1.8 K magnetization data shown before in Figure 3(b) fragments into a multitude of unique states at intermediate temperatures from 2–15 K. To better highlight the intricacy of the magnetic landscape near the potential 2/3 magnetization plateau, we collected higher density data over the limited ranges highlighted by the black dashed lines.
Figure 4(c,f) demonstrate the phase diagram for TbTi3Bi4 in the vicinity of the second metamagnetic transition. At least 5 unique phase pockets can be observed. Furthermore, while the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field. Data was also collected for the zero-field-cooled virgin curves, though they are qualitatively similar to Figure 4(a,b,c) and have been omitted for brevity. A series of line cuts (Figure 4(c,right)) through the phase diagram data (white dashed lines) helps to clarify the clear boundaries between the phase pockets. Considering the complexity of the magnetic landscape, neutron diffraction experiments will be invaluable to identify the types and natures of the individual magnetic phases.
Antiferromagnetic metals are not particularly common, particularly so for materials that can be readily exfoliated. While the low temperature magnetism of the rare-earth metals prevents facile applications within fields like spintronics, understanding how the magnetism interacts with the transport properties is key to understanding the connection between the kagome network (which dominates the electronic states near E) and the magnetism. Figure 5(a) shows a basic resistivity curve as a function of temperature under zero-field conditions. Our experiment passes current parallel to the [010] direction, and orients the field such that (oriented to induce the metamagnetic transitions). The crystal quality is good (RRR40) and the zero-field transport is indicative of a relatively high mobility metal. The magnetic transition at 20 K is evident in the resistivity, highlighted by a sharp cusp and subsequent drop. This is the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
We now apply a magnetic field to examine the magnetoresistance (MR) of TbTi3Bi4 in the same experimental setup. Figure 5(b) shows the symmetrized transverse magnetoresistance with and at 1.8 K. As one would expect from the magnetization results, we observe two sharp transitions in the magnetoresistance, one for each of the metamagnetic features. The magnetoresistance retains the qualitative field hysteresis of the magnetization data, both in magnitude and location. The first metamagnetic transition marks a sharp increase in the magnetoresistance of approximately 30–40%. The second metamagnetic transition marks a sharp decrease in the magnetoresistance to approximately 50%, which one may expect for a total polarization of the Tb moments. We find the sharp switching between positive and negative MR intriguing in TbTi3Bi4, particularly for a system which is already quite metallic (1Ohm-cm at 1.8 K).
Given the strong MR response in TbTi3Bi4, we were curious to investigate whether the complex landscape of magnetic transitions seen in the phase diagrams would influence the electronic transport. Figure 5(c) provides a series of magnetoresistance curves as a function of temperature from 2 K to 21 K in 1 K steps. For simplicity and clarity, we are only showing the data under increasing fields. We would like to highlight the behavior near the second transition of the symmetrized data (3 T). Immediately before the transition to negative magnetoresistance, we observe that the magnetoresistance cusp becomes a peak – then splits into two peaks – merge back together – and then vanish as the system approaches the Neel temperature. These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4(c). This further suggests that these are bulk effects corresponding to distinct types of magnetic order.
The sample geometry prevents us from measuring the Hall response where and . The in-plane moments would require thinning the sample perpendicular to the natural layers of the crystal, which typically causes the sample to flake and fall apart. However, we can readily orient the sample such that and to explore the Hall effect and magnetoresistance when the metamagnetic transitions are not induced. Figure 5(d) shows the antisymmetrized Hall resistivity for TbTi3Bi4 over two temperature regimes. The top figure shows the Hall response from 2 K to 18 K, and the bottom figure highlights temperatures above the magnetic transition temperature (20 K to 30 K). No subtractions or corrections have been performed besides antisymmetrization. All data sets exhibit a strongly nonlinear Hall response over all fields. A signal that looks deceptively like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature. Furthermore, at high temperatures we recover an additional inflection in the data. The signal is also relatively small, leading us to suspect that we may be seeing the effect of multiband transport. We can look to some prior work on NdTi3Bi4 for some guidance – since the electronic structure of these materials doesn’t appreciably change with the choice of rare-earth. Experimental ARPES data has observed a Dirac-like feature near the Fermi level[38, 37, 35, 36] and has further observed gapping of the Dirac-like bands as NdTi3Bi4 transitions into the magnetic ground state[35]. The sample exhibits strong positive magnetoresistance (200%) when and , in stark contrast to behavior when and .
Overall, our work into TbTi3Bi4 has revealed a unique antiferromagnetic, cleavable metal with an exceedingly complex landscape of magnetic phases. The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative magnetoresistance near the 2/3 magnetization plateau, and an unusual coupling between the magnetic phase pockets and the transport. We now turn to briefly investigate the magnetic properties of the Ln2-xTi6+xBi9 shurikagome metals and the influence of disorder on the Ln–Ln chains.
Shurikagome Ln | Occ | Occ | Tot Ln | Tot Ln | M () | M () | M () | T (K) | () | (K) |
Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.7 | -24 |
Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | -14 |
Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | -0.2 |
Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.6 | -0.9 |
Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 7.3 | 14 |
Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | 0 | 0 | 0.1 | N/A | N/A | N/A |
III.3 Properties of Ln2-xTi6+xBi9
Diffraction studies on the Ln2-xTi6+xBi9 compounds revealed several interesting observations. We first attempted to solve the structure using fully occupied Ln sites (forming fully occupied chains). While R values were actually quite favorable, there existed large amounts of excess electron density around the Ln04 and Ln05 sites. The residual Q-peaks can be eliminated by allowing the occupancy of the Ln sites to vary, which results in partially occupied Ln chains. However, compositions extracted from single-crystal X-ray diffraction (SCXRD) using the assumption of pure (although partially occupied) Ln sites are in strong disagreement with electron dispersive spectroscopy (EDS) results. Bluntly, there is not enough Ti in the unit cell derived from SCXRD. This can be remedied by allowing antisite disorder on the Ln sites, allowing the composition of the Ln to vary, but restraining it such that the site remains fully occupied (e.g. Ln04 + Ti04 = 1). Even this appears to slightly underestimate the amount of Ti, and EDS results suggest that some Ti may be substituting elsewhere as well. This is perhaps expected, as the published Ti–Bi phase diagram indicates a full solid solution over the range of 50:50 Ti:Bi to 35:65 Ti:Bi.[41]
However, the SCXRD data did not independently suggest any additional site substitutions, and the constraints mentioned above yielded quite favorable statistics. Figure 6(a) highlights the cell volume for the Ln2-xTi6+xBi9 shurikagome metals as a function of the 8-coordinate Shannon radius. Despite the variable rare-earth occupancy, the compounds still adhere to a nearly linear expansion in cell volume, as one would expect of a rare-earth series. These results are briefly summarized in Table 1. While CIF files are provided in the ESI, here we focus on the influence of the Ln atom and the disorder in crystals of Ln2-xTi6+xBi9. The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln04 and Ln05. As mentioned before, the Ti occupancy on those sites is constrained such that they add to 1. The sum of the SCXRD-determined occupancies is provided in the column labeled “Tot Ln.” This can be directly compared with the composition extracted from EDS measurements in the next column.
SCXRD appears to consistently calculate higher concentrations of Ln as compared to EDS measurements, over approximating by about 15%. Some interesting trends appear in the overall data. Both the Ln04 and Ln05 site occupancies decrease as we move to smaller Ln atoms. This effect is substantially more pronounced on the Ln05 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9. We performed some exploratory syntheses around Tb2-xTi6+xBi9 in an attempt to fully saturate the Ln site, though only small changes in the rare-earth composition were noted.
Considering the high disorder in crystals of Ln2-xTi6+xBi9, particularly within the Ln sublattices, it is not clear what magnetic response to expect. Naively, the most likely scenario was freezing into a spin-glass, or a complete loss of a well-defined magnetic ground state in the non-Kramers ions. We now turn to a brief overview of the magnetic properties of the Ln2-xTi6+xBi9 compounds.
The compounds are tetragonal (P4/nmm) yielding two directions of interest, (parallel to chain) and (perpendicular to chain). We did not consider rotations with in the ab plane for this work. All compounds exhibit weak anisotropy with the exception of Tm. When , Tm2-xTi6+xBi9 exhibits an order of magnitude stronger response. For Tb, Dy, and Ho, data collected with exhibits a cusp in the magnetization, strangely resembling an antiferromagnetic transition. For Er and Tm, the data with exhibits a broad peak instead. All other combinations of elements and directions (e.g. Tb with are featureless, although the magnitude of the magnetization remains comparable to the other direction). These results are summarized in Figure 6(a) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction. Please note that the magnetization results are shown in arbitrary units for ease of comparison. Full data sets are provided in the ESI.
Note that the data presented in Figure 6(a) includes both zero-field-cooled (ZFC) and field-cooled (FC) data. The lack of ZFC/FC splitting is surprising, as we may reasonably expect the strong disorder to result in a glassy magnetic state. Furthermore, AC susceptibility measurements on Tb2-xTi6+xBi9 indicate no frequency dependence. However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits a strong frequency response. It is striking that the features follow the temperature scale expected of the latter half of the rare-earth row – Tb possesses the highest ordering temperature, and Tm possesses the lowest.
Figure 6(b) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K. Again we have opted to show only one direction (the same direction as in Figure 6(a)) for simplicity. All compounds near saturation by 7 T, though Er2-xTi6+xBi9 has not saturated, owing to an unusual inflection in the data near 4 T. Regardless, we can extract the approximate saturation magnetization for all compounds, which has been summarized in Table 1 (column M). Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2-xTi6+xBi9. This will underestimate the true saturation magnetization, as we will be normalizing to 2 rare-earth atoms per formula unit instead of (2-x). However, we can compare these results with the expected saturation magnetization from XRD and EDS results (M and M). Comparing directly, we find that the experimental results actually agree quite well with the SCXRD results. Some other parameters of interest (e.g. and ) have also been added to Table 1. Due to the disorder and variable number of Ln sites, we consider these values as first approximations.
Despite the high levels of disorder, some samples of Ln2-xTi6+xBi9 (e.g. Tb) appear to show magnetization consistent with long-range order. This is highly unexpected, and considering the lack of ZFC/FC splitting and no clear frequency dependence, more detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed.
IV Conclusions
We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, a kagome metal belonging to the AM3X4 family of materials. We described the rational for the formation of TbTi3Bi4, including the increased competition with another class of materials, the Ln2-xTi6+xBi9 shurikagome metals for small rare-earth species (Ln: Tb–Lu). Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases surrounding multiple metamagnetic transitions. Transport measurements highlight the impact the rare-earth magnetism has on the kagome-dominated transport. Upon entering the magnetization plateau, the system sharply switches between positive (30-40%) and negative (-55%) magnetoresistance in a system that is already quite metallic (1 Ohm-cm). We have also briefly examined the properties of the Ln2-xTi6+xBi9 shurikagome metals. Despite considerable disorder on the rare-earth sublattices, some systems exhibit magnetization responses more indicative of long-range order. Together, these results not only provide insight into a new kagome lattice and an unusual family of disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
V Acknowledgements
The work of B.R.O (synthesis, data collection, writing), H.Z. (analysis), K.G. (He3 data collection), A.F.M. (He3 data collection), and M.A.M. (editing), was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. We thank the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source for use of their Photonics Laue Camera System.